The owner of a delicatessen mixed coffee that sold for $4.50 per pound with coffee that sold for $3.00 per pound. How many pounds of each were used to make 20 pounds of a blend that would sell for $3.90 per pound?
Answer by checkley77(12844)

4.5X+3(20-X)=3.9*20
4.5X+60-3X=78
1.5X=78-60
1.5X=18
X=18/1.5
X=12 POUNDS OF $4.50 COFFEE IS USED.
20-12=8 POUNDS OF $3.00 COFFEE IS USED.
PROOF:
4.5*12+3*8=78
54+24=78
78=78